OpenStudy (anonymous):

s=1/2gt^2 solve for t

jimthompson5910 (jim_thompson5910):

\[\Large s = \frac{1}{2}gt^2\] \[\Large 2s = gt^2\] \[\Large \frac{2s}{g} = t^2\] \[\Large t^2 = \frac{2s}{g}\] With me so far?